# Integral cohomology rings of weighted Grassmann orbifolds and rigidity properties

### 11 Mar 2024 v2

@@ 5.2. Structure constants for the equivariant cohomology of $\mbox{Gr}_{\bf b}(k,n)$
[…]
Note that there exists an isomorphism
$\phi^{*}:H^{*}(BT_{\bf{b}};\mathbb{Q})\to H^{*}(BT^{n};\mathbb{Q})$
defined by $\phi^{*}(x_{i})=y_{i}$. Observe that $\phi^{*}(f)=f|_{z=0}\in\mathbb{Q}[y_{1},\dots,y_{n}]$ for a polynomial $f\in\mathbb{Q}[x_{1},\dots,x_{n}]$.
Proposition 5.9.
The $T_{\bf{b}}$-equivariant cohomology of $\mbox{Gr}_{\bf{b}}(k,n)$ with rational coefficient is given by
$H_{T_{\bf{b}}}^{*}(\mbox{Gr}_{\bf{b}}(k,n);\mathbb{Q})=\Big{\{}\alpha\in\bigoplus_{j=0}^{m}\mathbb{Q}[x_{1},\dots,x_{n}]~{}\big{|}~{}(X_{\lambda^{i}}-d_{ij}X_{\lambda^{j}})~{}\mbox{divides}~{}(\alpha_{j}-\alpha_{i})\text{ if }\lambda^{i}\in R(\lambda^{j})\Big{\}}.$
Here $X_{\lambda}=\sum_{i=1}^{k}x_{\lambda_{i}}$ for every Schubert symbol $\lambda=(\lambda_{1},\dots,\lambda_{k})$.
@@
Since $\pi^{*}$ is a ring isomorphism, $\{p\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$ is a basis for $H^{*}_{T^{n+1}}(Pl(k,n);\mathbb{Q})$. Note that $f\in\mathbb{Q}[y_{1},\dots,y_{n},z]$ then $f|_{z=0}\in\mathbb{Q}[y_{1},\dots,y_{n}]$.
@@
${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}=(\pi^{*}_{{\bf{b}}})^{-1}(p\widetilde{\mathbb{S}}_{\lambda^{i}}|_{z=0}).$
@@
$\widetilde{\mathbb{S}}_{\lambda^{i}}|_{\lambda^{i}}=\prod_{\lambda^{\ell}\in R(\lambda^{i})}(Y_{\lambda^{\ell}}-Y_{\lambda^{i}})$. Thus
@@
Note that if $f\in\frac{\mathbb{Q}[y_{1},\dots,y_{n},z]}{(Y_{\lambda^{i}}+z)}$ then $f|_{z=0}\in\frac{\mathbb{Q}[y_{1},\dots,y_{n}]}{(Y_{\lambda^{i}})}$. Thus
$(p\widetilde{\mathbb{S}}_{\lambda^{i}}|_{\lambda^{i}})|_{z=0}=\prod_{\lambda^{\ell}\in R(\lambda^{i})}Y_{\lambda^{\ell}}\in\frac{\mathbb{Q}[y_{1},\dots,y_{n}]}{(Y_{\lambda^{i}})}.$
[…]
Since $(\pi_{b})^{*}$ is an isomorphism, $p\widetilde{\mathbb{S}}_{\lambda^{i}}|_{\lambda^{j}}\neq 0\implies\lambda^{j}\succeq\lambda^{i}$. Now the proof follows from the uniqueness of ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}$. ∎
The set $\{{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$ forms a basis for the equivariant cohomology ring $H_{T^{n}}^{*}(\mbox{Gr}_{\bf{b}}(k,n);\mathbb{Q})$ as well. To compute the equivariant structure constants with respect to the basis $\{{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$, we utilize Lemma 5.7, along with the isomorphisms $\pi^{*}$ and $\pi_{\bf{b}}^{*}$, as well as Lemma 5.10. A priory, these constants are expected to be elements of $\mathbb{Q}[y_{1},y_{2},\dots,y_{n}]$. However, we proceed to establish that all the equivariant structure constants lie in $\mathbb{Z}[y_{1},y_{2},\dots,y_{n}]$ for a divisive weighted Grassmann orbifold.
Using Lemma 5.7 and (5.3) we have
(5.4) $p\widetilde{\mathbb{S}}_{\lambda^{i}}p\widetilde{\mathbb{S}}_{\lambda^{j}}=\sum_{\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\sum_{P:\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}}\sigma_{r}wt(P)p\widetilde{\mathbb{S}}_{\lambda^{q}},$
where
$\sigma_{r}wt(P)=\prod_{s=1}^{|P|}\sigma_{r}wt(p_{s})$
and $p_{s}$ ranges over the equivariant pieces on $P$ such that $\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}$. Observe that $|P|=d(\lambda^{i})+d(\lambda^{j})-d(\lambda^{q})$ for a puzzle $P$ such that $\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}$. Note that $\sigma_{r}wt(p_{s})=Y_{\lambda^{e}}-Y_{\lambda^{f}}$ for some $\lambda^{e}\prec\lambda^{f}$. We define
(5.5) ${\bf{b}}(p_{s}):=b_{e}-b_{f}.$
@@
Now, $\pi^{*}$ is a ring isomorphism and $p\widetilde{\mathbb{S}}_{\lambda^{1}}=\pi^{*}(\widetilde{\mathbb{S}}_{\lambda^{1}}).$ Thus
@@
Recall that $H^{*}_{T^{n+1}}(Pl(k,n);\mathbb{Q})$ is also a $\mathbb{Q}[x_{1},\dots,x_{n}]$ module and $(\pi_{\bf{b}})^{*}$ is a ring isomorphism over $\mathbb{Q}[x_{1},\dots,x_{n}]$ module. Next, we relate two different module structures on $H^{*}_{T^{n+1}}(Pl(k,n);\mathbb{Q})$. Note that $X_{\lambda^{i}}=Y_{\lambda^{i}}-\frac{b_{i}-a}{a}z$ for all $i=0,1,\dots,m$. Thus
$Y_{\lambda^{e}}-Y_{\lambda^{f}}=X_{\lambda^{e}}-X_{\lambda^{f}}+\frac{b_{e}-b_{f}}{a}z=X_{\lambda^{e}}-X_{\lambda^{f}}+\frac{b_{e}-b_{f}}{b_{0}}(Y_{\lambda^{0}}-X_{\lambda^{0}}+z).$
Using Proposition 5.11, define
$K:=X_{\lambda^{e}}-X_{\lambda^{f}}-\frac{b_{e}-b_{f}}{b_{0}}X_{\lambda^{0}}+\frac{b_{e}-b_{f}}{b_{0}}p\widetilde{\mathbb{S}}_{\lambda^{1}}\in H^{*}_{T^{n+1}}(Pl(k,n);\mathbb{Q}).$
Thus $K\cdot p\widetilde{\mathbb{S}}_{\lambda^{q}}\in H^{*}_{T^{n+1}}(Pl(k,n);\mathbb{Q})$ and
$\sigma_{r}wt(p_{s})\cdot p\widetilde{\mathbb{S}}_{\lambda^{q}}=K\cdot p\widetilde{\mathbb{S}}_{\lambda^{q}}\text{ for all }q=0,1,\dots,m.$
Then evaluating at $z=0$ on both sides we get the following relation in $H^{*}_{T^{n+1}}(Pl(k,n);\mathbb{Q})$.
$\sigma_{r}wt(p_{s})(p\widetilde{\mathbb{S}}_{\lambda^{q}}|_{z=0})={\bf{b}}wt(p_{s})(p\widetilde{\mathbb{S}}_{\lambda^{q}}|_{z=0})+\frac{{\bf{b}}(p_{s})}{b_{0}}(p\widetilde{\mathbb{S}}_{\lambda^{1}}|_{z=0})(p\widetilde{\mathbb{S}}_{\lambda^{q}}|_{z=0}),$
where ${\bf{b}}wt(p_{s})=(Y_{\lambda^{e}}-Y_{\lambda^{f}}-\frac{{\bf{b}}(p_{s})}{b_{0}}Y_{\lambda^{0}})$. Thus
(5.6) $\prod_{s=1}^{|P|}\sigma_{r}wt(p_{s})(p\widetilde{\mathbb{S}}_{\lambda^{q}}|_{z=0})$ $=\prod_{s=1}^{|P|}({\bf{b}}wt(p_{s})+\frac{{\bf{b}}(p_{s})}{b_{0}}(p\widetilde{\mathbb{S}}_{\lambda^{1}}|_{z=0}))(p\widetilde{\mathbb{S}}_{\lambda^{q}}|_{z=0})$
$=\sum_{s=0}^{|P|}a^{(P)}_{s}(p\widetilde{\mathbb{S}}_{\lambda^{1}}|_{z=0})^{s}(p\widetilde{\mathbb{S}}_{\lambda^{q}}|_{z=0}),$
where $a^{(P)}_{s}$ can be calculated using the following remark.
@@
If $\prod_{s=1}^{t}(a_{s}+b_{s}x)=\sum_{s=0}^{t}\alpha_{s}x^{s}$ then
$\alpha_{s}=\sum_{1\leq i_{1}<i_{2}\dots<i_{s}\leq t}b_{i_{1}}b_{i_{2}}\dots b_{i_{s}}\frac{a_{1}a_{2}\dots a_{t}}{a_{i_{1}}a_{i_{2}}\dots a_{i_{s}}}.$
Then using (5.4), (5.6) and Lemma 5.10, we have
${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{j}}$ $=\sum_{\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\sum_{P:\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}}\sum_{s=0}^{|P|}a^{(P)}_{s}({\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{1}})^{s}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{q}}.$
We have the weighted Pieri rule for the multiplication with the equivariant Schubert class ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{1}}$. For any two Schubert symbol $\lambda^{i}$ and $\lambda^{j}$ we denote $\lambda^{j}\to\lambda^{i}$ if $d(\lambda^{j})=d(\lambda^{i})+1$ and $\lambda^{i}\preceq\lambda^{j}$.
Proposition 5.13 (Equivariant Pieri rule).
[7, Proposition 5.7]
${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{1}}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}=(Y_{\lambda^{0}}-\frac{b_{0}}{b_{i}}Y_{\lambda^{i}}){\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}+\sum_{\lambda^{j}:\lambda^{j}\to\lambda^{i}}\frac{b_{0}}{b_{i}}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{j}}.$
For $s\geq 1$, let ${\bf{b}}\tilde{\mathscr{C}}_{q~{}g^{s}}^{\ell}\in\mathbb{Z}[y_{1},y_{2},\dots,y_{n}]$ which satisfies
(5.7) $({\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{1}})^{s}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{q}}=\sum_{\lambda^{\ell}\succeq\lambda^{q}}{\bf{b}}\widetilde{\mathscr{C}}_{q~{}g^{s}}^{\ell}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{\ell}}.$
Now evaluating the Proposition 5.13 $s$ times for $s\geq 1$, we can compute the equivariant structure constants ${\bf{b}}\tilde{\mathscr{C}}_{q~{}g^{s}}^{\ell}$ by the following:
${\bf{b}}\widetilde{\mathscr{C}}_{q~{}g^{s}}^{\ell}=\sum_{\begin{subarray}{c}\lambda^{\ell}=\lambda^{\ell_{0}}\to\lambda^{\ell_{1}}\to\\ \dots\to\lambda^{\ell_{r}}=\lambda^{q}\end{subarray}}\frac{(b_{0})^{r}}{b_{\ell_{1}}b_{\ell_{2}}\dots b_{\ell_{r}}}\sum_{J}(\prod_{t=0}^{r}(Y_{\lambda^{0}}-\frac{b_{0}}{b_{\ell_{t}}}Y_{\lambda^{\ell_{t}}})^{j_{t}}),$
where $J$ runs over all sequences $(j_{0},j_{1},\dots,j_{r})$ of non negative integers satisfying $j_{0}+j_{1}+\dots+j_{r}=s-r$. Note that $r=d(\lambda^{\ell})-d(\lambda^{q})$. Thus ${\bf{b}}\widetilde{\mathscr{C}}_{q~{}g^{s}}^{\ell}=0$ if $s<d(\lambda^{\ell})-d(\lambda^{q})$.
@@
${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}=\sum_{\lambda^{\ell}\succeq\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\sum_{P:\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}}\sum_{s=0}^{|P|}a^{(P)}_{s}{\bf{b}}\widetilde{\mathscr{C}}_{q~{}g^{s}}^{\ell}.$
@@
we have
${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{j}}$ $=\sum_{\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\left(\sum_{P:\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}}\sum_{s=0}^{|P|}a^{(P)}_{s}({\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{1}})^{s}\right){\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{q}}$
(5.8) $=\sum_{\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\sum_{P:\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}}\sum_{s=0}^{|P|}a^{(P)}_{s}\left(\sum_{\lambda^{\ell}\succeq\lambda^{q}}{\bf{b}}\widetilde{\mathscr{C}}_{q~{}g^{s}}^{\ell}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{\ell}}\right)$
$=\sum_{\lambda^{\ell}}\left(\sum_{\lambda^{\ell}\succeq\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\sum_{P:\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}}\sum_{s=0}^{|P|}a^{(P)}_{s}{\bf{b}}\widetilde{\mathscr{C}}_{q~{}g^{s}}^{\ell}\right){\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{\ell}}$
Thus to calculate the equivariant structure constant ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ first we need to find and sum over all $\lambda^{q}$ which satisfies ${\lambda^{\ell}\succeq\lambda^{q}\succeq\lambda^{i},\lambda^{j}}$. For each such $\lambda^{q}$ we need to sum over all equivariant puzzle $P$ such that $\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}$. The number of equivariant puzzle pieces of $P$ is $|P|=d(\lambda^{i})+d(\lambda^{j})-d(\lambda^{q})$.
Now we need to prove that ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}\in\mathbb{Z}[y_{1},y_{2},\dots,y_{n}]$ for all $i,j$ and $\ell$. The argument follows by using ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}|_{\lambda^{i}}$ is a primitive polynomial in $\mathbb{Z}[y_{1},y_{2},\dots,y_{n}]$ and the upper triangularity of the equivariant Schubert classes ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}$. We recall a polynomial in $\mathbb{Z}[y_{1},y_{2},\dots,y_{n}]$ is said to be primitive if the gcd of all coefficients of the monomials is one.

### 09 Mar 2026 v3

@@ 4. Torsion’s in the integral cohomology of weighted Grassmann orbifold
$\text{ if }|\lambda^{i}\cap\lambda^{j}|=k-1\Big\}.$
@@
Since $\pi^{*}$ is an algebra isomorphism, $\{P\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$ is a basis for $H^{*}_{T^{n+1}}(Pl(k,n);\mathbb{Q})$ as a $\mathbb{Q}[y_{1},\dots,y_{n}]$-algebra. Moreover, $(\pi^{*}_{{\bf{b}}})^{-1}(P\widetilde{\mathbb{S}}_{\lambda^{i}})$ forms a basis of $H_{T_{\bf{b}}}^{*}(\mbox{Gr}_{\bf{b}}(k,n);\mathbb{Q})$ as a $\mathbb{Q}[x_{1},\dots,x_{n}]$-algebra. The next Lemma follows from the the diagram in Figure 1 and the uniqueness of the equivariant Schubert class ${\bf b}\mathbb{S}_{\lambda^{i}}$.
@@
${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}=\phi^{*}\circ(\pi^{*}_{{\bf{b}}})^{-1}(P\widetilde{\mathbb{S}}_{\lambda^{i}}).$
@@
From Definition 5.2, it follows $\widetilde{\mathbb{S}}_{\lambda^{i}}|_{\lambda^{i}}=\prod_{\lambda^{\ell}\in R(\lambda^{i})}(Y_{\lambda^{\ell}}-Y_{\lambda^{i}})$. Thus
@@
Since $x_{r}=y_{r}-\frac{w_{r}}{a}z$, we have $X_{\lambda^{i}}=Y_{\lambda^{i}}-(\frac{b_{i}-a}{a})z$. Thus
$\prod_{\lambda^{\ell}\in R(\lambda^{i})}(X_{\lambda^{\ell}}-\dfrac{b_{\ell}}{b_{i}}X_{\lambda^{i}})=\prod_{\lambda^{\ell}\in R(\lambda^{i})}(Y_{\lambda^{\ell}}+z)-\dfrac{b_{\ell}}{b_{i}}(Y_{\lambda^{i}}+z)=\prod_{\lambda^{\ell}\in R(\lambda^{i})}(Y_{\lambda^{\ell}}+z)\in{\mathbb{Q}[y_{1},\dots,y_{n}]}/{(Y_{\lambda^{i}}+z)}.$
$\text{ Also, }\phi^{*}\Big(\prod_{\lambda^{\ell}\in R(\lambda^{i})}(X_{\lambda^{\ell}}-\dfrac{b_{\ell}}{b_{i}}X_{\lambda^{i}})\Big)=\prod_{\lambda^{\ell}\in R(\lambda^{i})}(Y_{\lambda^{\ell}}-\dfrac{b_{\ell}}{b_{i}}Y_{\lambda^{i}})={\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}|_{\lambda^{i}}.$
Since $\pi^{*}$ is an isomorphism, $P\widetilde{\mathbb{S}}_{\lambda^{i}}|_{\lambda^{j}}\neq 0\implies\lambda^{j}\succeq\lambda^{i}$. Similarly, $\phi^{*}\circ(\pi^{*}_{{\bf{b}}})^{-1}(P\widetilde{\mathbb{S}}_{\lambda^{i}})$ satisfies the same property as $P\widetilde{\mathbb{S}}_{\lambda^{i}}$. Thus $P\widetilde{\mathbb{S}}_{\lambda^{i}}$ satisfies all the assumptions of Definition 5.2. Now the proof follows from the uniqueness of ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}$. ∎
For the sake of computing the equivariant structure constants ${\bf{b}}\widetilde{\mathscr{C}}_{i~j}^{\ell}$, we view the product $U_{R}{\cdot}P\widetilde{\mathbb{S}}_{\lambda^{q}}$ in $H_{T^{n+1}}^{*}(Pl(k,n))$ as a $\mathbb{Q}[x_{1},\dots x_{n}]$ algebra. Note that
(5.5) $U_{R}=\prod_{r\in R}(y_{r}-y_{r+1})=\prod_{r\in R}\Big((x_{r}-x_{r+1})+(w_{r}-w_{r+1})\frac{z}{a}\Big)=\sum_{s=0}^{|R|}L_{s,R}(\frac{z}{a})^{s}.$
The coefficients $L_{s,R}$ can be computed by the following, where we consider $S$ as a sub collection of $R$ of cardinality $s$
(5.6) $L_{s,R}=\sum_{S:~|S|=s}\Big(\prod_{r\in S}(w_{r}-w_{r+1})\prod_{r\in R\setminus S}(x_{r}-x_{r+1})\Big)$
Remark 5.7.
If $s=|R|$, then $L_{s,R}=\prod_{r\in R}(w_{r}-w_{r+1})$.
For, $k<n$, the smallest Schubert symbol $\lambda^{0}$ is given by $\lambda^{0}=(1,2,\dots,k)$. Moreover the Schubert symbol $\lambda^{1}$ is given by $\lambda^{1}=(1,2,\dots,k-1,k+1)$.
@@
Since, $P\widetilde{\mathbb{S}}_{\lambda^{1}}=\pi^{*}(\widetilde{\mathbb{S}}_{\lambda^{1}})$ we have
@@
For any two Schubert symbol $\lambda^{i}$ and $\lambda^{j}$, we denote $\lambda^{j}\to\lambda^{i}$ if $d(\lambda^{j})=d(\lambda^{i})+1$ and $\lambda^{i}\preceq\lambda^{j}$. Following [16, Proposition 2], and Proposition 5.8, the Pieri rule in $H_{T^{n+1}}^{*}(Pl(k,n))$ as a $\mathbb{Q}[y_{1},\dots,y_{n}]$ algebra is given by
$(Y_{\lambda^{0}}+z)P\widetilde{\mathbb{S}}_{\lambda^{q}}$ $=(Y_{\lambda^{0}}-Y_{\lambda^{q}})P\widetilde{\mathbb{S}}_{\lambda^{q}}+\sum_{\lambda^{\ell}:~\lambda^{\ell}\to\lambda^{q}}P\widetilde{\mathbb{S}}_{\lambda^{\ell}}.$
(5.7) $\implies z{\cdot}P\widetilde{\mathbb{S}}_{\lambda^{q}}$ $=-Y_{\lambda^{q}}P\widetilde{\mathbb{S}}_{\lambda^{q}}+\sum_{\lambda^{\ell}:~\lambda^{\ell}\to\lambda^{q}}P\widetilde{\mathbb{S}}_{\lambda^{\ell}}.$
The ring $H_{T^{n+1}}^{*}(Pl(k,n))$ can also be considered as a $\mathbb{Q}[x_{1},\dots,x_{n}]$ algebra. In order to describe the above product as a $\mathbb{Q}[x_{1},\dots,x_{n}]$ algebra we substitute $y_{r}=x_{r}+\frac{w_{r}}{a}z$ for all $r=1,2,\dots,n$ in the RHS of (5). Thus
$z{\cdot}P\widetilde{\mathbb{S}}_{\lambda^{q}}=-(X_{\lambda^{q}}+\frac{b_{q}-a}{a}z)P\widetilde{\mathbb{S}}_{\lambda^{q}}+\sum_{\lambda^{\ell}:~\lambda^{\ell}\to\lambda^{q}}P\widetilde{\mathbb{S}}_{\lambda^{\ell}}$
(5.8) $\implies\frac{z}{a}{\cdot}P\widetilde{\mathbb{S}}_{\lambda^{q}}=-\frac{1}{b_{q}}X_{\lambda^{q}}P\widetilde{\mathbb{S}}_{\lambda^{q}}+\frac{1}{b_{q}}\sum_{\lambda^{\ell}:~\lambda^{\ell}\to\lambda^{q}}P\widetilde{\mathbb{S}}_{\lambda^{\ell}}.$
@@
The explicit relation between the $\mathbb{Q}[x_{1},\dots,x_{n}]$ algebra structures and $\mathbb{Q}[y_{1},\dots,y_{n}]$ algebra structures on $H_{T^{n+1}}^{*}(Pl(k,n))$ follows from (5). More explicitly, for $1\leq r\leq n$
$y_{r}P\widetilde{\mathbb{S}}_{\lambda^{q}}=(x_{r}+\frac{w_{r}}{a}z)P\widetilde{\mathbb{S}}_{\lambda^{q}}=x_{r}P\widetilde{\mathbb{S}}_{\lambda^{q}}-\frac{w_{r}}{b_{q}}X_{\lambda^{q}}P\widetilde{\mathbb{S}}_{\lambda^{q}}+\frac{w_{r}}{b_{q}}\sum_{\lambda^{\ell}:~\lambda^{\ell}\to\lambda^{q}}P\widetilde{\mathbb{S}}_{\lambda^{\ell}}.$
For $s\geq 1$, repeated evaluation of the product in (5) leads to the following formula.
$(\frac{z}{a})^{s}P\widetilde{\mathbb{S}}_{\lambda^{q}}=\sum_{\lambda^{\ell}\succeq\lambda^{q}}{\bf{b}}\widetilde{{C}}_{q~g^{s}}^{\ell}P\widetilde{\mathbb{S}}_{\lambda^{\ell}}.$
The coefficients ${\bf{b}}\widetilde{{C}}_{q~g^{s}}^{\ell}$ is given by the following:
(5.9) ${\bf{b}}\widetilde{{C}}_{q~g^{s}}^{\ell}=\sum_{\begin{subarray}{c}\lambda^{\ell}=\lambda^{\ell_{0}}\to\lambda^{\ell_{1}}\to\\ \dots\to\lambda^{\ell_{r}}=\lambda^{q}\end{subarray}}\frac{1}{b_{\ell_{1}}b_{\ell_{2}}\dots b_{\ell_{r}}}\sum_{J}(\prod_{t=0}^{r}(-\frac{1}{b_{\ell_{t}}}X_{\lambda^{\ell_{t}}})^{j_{t}}),$
where $J$ runs over all sequences $(j_{0},j_{1},\dots,j_{r})$ of non negative integers satisfying $j_{0}+j_{1}+\dots+j_{r}=s-r$. Note that $r=d(\lambda^{\ell})-d(\lambda^{q})$. Thus ${\bf{b}}\widetilde{{C}}_{q~g^{s}}^{\ell}=0$ if $d(\lambda^{\ell})>s+d(\lambda^{q})$.
Remark 5.10.
If $d(\lambda^{\ell})=d(\lambda^{q})+s$ then
${\bf{b}}\widetilde{{C}}_{q~g^{s}}^{\ell}=\sum_{\begin{subarray}{c}\lambda^{\ell}=\lambda^{\ell_{0}}\to\lambda^{\ell_{1}}\to\\ \dots\to\lambda^{\ell_{s}}=\lambda^{q}\end{subarray}}\frac{1}{b_{\ell_{1}}b_{\ell_{2}}\dots b_{\ell_{s}}}.$
We define $\mathcal{L}_{s,R}:=\phi^{*}(L_{s,R})$ and ${\bf{b}}\widetilde{\mathscr{C}}_{q~g^{s}}^{\ell}:=\phi^{*}({\bf{b}}\widetilde{{C}}_{q~g^{s}}^{\ell})$. In other words, if we substitute $x_{i}$ by $y_{i}$ in $L_{s,R}$ and ${\bf{b}}\widetilde{{C}}_{q~g^{s}}^{\ell}$ we get $\mathcal{L}_{s,R}$ and ${\bf{b}}\widetilde{\mathscr{C}}_{q~g^{s}}^{\ell}$ respectively.
@@
${\bf{b}}\widetilde{\mathscr{C}}_{i~j}^{\ell}=\sum_{\lambda^{\ell}\succeq\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\sum_{R}\sum_{s=0}^{|R|}K(i,j,q,R)\mathcal{L}_{s,R}{\bf{b}}\widetilde{\mathscr{C}}_{q~g^{s}}^{\ell}.$
@@
Using Lemma 5.6 in (5) together with (5.5) we have
${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{j}}$ $=\sum_{|R|=d(\lambda^{i})+d(\lambda^{j})-d(\lambda^{q})}K(i,j,q,R)\sum_{s=0}^{|R|}\mathcal{L}_{s,R}(\frac{z}{a})^{s}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{q}}$
(5.10) $=\sum_{|R|=d(\lambda^{i})+d(\lambda^{j})-d(\lambda^{q})}K(i,j,q,R)\sum_{s=0}^{|R|}\mathcal{L}_{s,R}\sum_{\lambda^{\ell}\succeq\lambda^{q}}{\bf{b}}\widetilde{\mathscr{C}}_{q~g^{s}}^{\ell}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{\ell}}$
$=\sum_{\lambda^{\ell}}\left(\sum_{\lambda^{\ell}\succeq\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\sum_{R}\sum_{s=0}^{|R|}K(i,j,q,R)\mathcal{L}_{s,R}{\bf{b}}\widetilde{\mathscr{C}}_{q~g^{s}}^{\ell}\right){\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{\ell}}$
Corollary 5.12 (Equivariant Pieri rule).
[7, Proposition 5.7]
${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{1}}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}=(Y_{\lambda^{0}}-\frac{b_{0}}{b_{i}}Y_{\lambda^{i}}){\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}+\frac{b_{0}}{b_{i}}\sum_{\lambda^{j}:\lambda^{j}\rightarrow\lambda^{i}}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{j}}.$
Proof.
Following [16, Proposition 2], and the isomorphism $\pi^{*}$ the Pieri rule in $H_{T^{n+1}}^{*}(Pl(k,n))$ is given by
(5.11) $P\widetilde{\mathbb{S}}_{\lambda^{1}}P\widetilde{\mathbb{S}}_{\lambda^{i}}$ $=(Y_{\lambda^{0}}-Y_{\lambda^{i}})P\widetilde{\mathbb{S}}_{\lambda^{i}}+\sum_{\lambda^{j}:\lambda^{j}\rightarrow\lambda^{i}}P\widetilde{\mathbb{S}}_{\lambda^{j}}$
To describe the product as a $\mathbb{Q}[x_{1},\dots,x_{n}]$-algebra we substitute $Y_{\lambda^{i}}=X_{\lambda^{i}}+\frac{b_{i}-a}{a}z$. Thus
$(Y_{\lambda^{0}}-Y_{\lambda^{i}})P\widetilde{\mathbb{S}}_{\lambda^{i}}$ $=(X_{\lambda^{0}}-X_{\lambda^{i}}+(b_{0}-b_{i})\frac{z}{a})P\widetilde{\mathbb{S}}_{\lambda^{i}}$
[…]
$=(X_{\lambda^{0}}-\frac{b_{0}}{b_{i}}X_{\lambda^{i}})P\widetilde{\mathbb{S}}_{\lambda^{i}}+\frac{b_{0}-b_{i}}{b_{i}}\sum_{\lambda^{j}:\lambda^{j}\rightarrow\lambda^{i}}P\widetilde{\mathbb{S}}_{\lambda^{j}}$
Then the proof follows from (5.11) using Lemma 5.6. ∎
Next, we prove that ${\bf{b}}\widetilde{\mathscr{C}}_{i~j}^{\ell}\in\mathbb{Z}[y_{1},y_{2},\dots,y_{n}]$ for all $i,j$ and $\ell$. We call a polynomial in $\mathbb{Z}[y_{1},y_{2},\dots,y_{n}]$ is primitive if the gcd of all coefficients of the monomials is one.